# SongGen: A Single Stage Auto-regressive Transformer for Text-to-Song Generation

### v1

@@ Abstract
Text-to-song generation, the task of creating vocals and accompaniment from textual inputs, poses significant challenges due to domain complexity and data scarcity. Existing approaches often employ multi-stage generation procedures, resulting in cumbersome training and inference pipelines. In this paper, we propose SongGen, a fully open-source, single-stage auto-regressive transformer designed for controllable song generation. The proposed model facilitates fine-grained control over diverse musical attributes, including lyrics and textual descriptions of instrumentation, genre, mood, and timbre, while also offering an optional three-second reference clip for voice cloning. Within a unified auto-regressive framework, SongGen supports two output modes: mixed mode, which generates a mixture of vocals and accompaniment directly, and dual-track mode, which synthesizes them separately for greater flexibility in downstream applications. We explore diverse token pattern strategies for each mode, leading to notable improvements and valuable insights. Furthermore, we design an automated data preprocessing pipeline with effective quality control. To foster community engagement and future research, we will release our model weights, training code, annotated data, and preprocessing pipeline. The generated samples are showcased on our project page at https://liuzh-19.github.io/SongGen/, and the code will be available at https://github.com/LiuZH-19/SongGen.
@@ 1 Introduction
Figure 1: Traditional methods often rely on multi-stage processes, making pipelines inflexible and complex. SongGen simplifies this with a single-stage auto-regressive transformer that supports both mixed mode and dual-track mode song generation.
Song generation presents greater complexity than speech or instrumental music generation (Lyth & King, 2024; Chen et al., 2024b; Liu et al., 2024; Copet et al., 2024), and the scarcity of open-source data further limits research in this area. Current approaches (Hong et al., 2024; Li et al., 2024a) tackle this challenge by separating vocals and accompaniment into distinct tracks, relying on multi-stage generation processes. As illustrated in Figure 1, these models first generate the vocal track from lyrics, then produce the accompaniment using natural language prompts alongside the generated vocals. However, multi-stage generation results in cumbersome training and inference pipelines while lacking unified control over both vocals and accompaniment. To improve pipeline simplicity and control flexibility, an important question arises: Is it possible for a single-stage model to achieve effective text-to-song generation?
@@
However, due to the sophisticated relationship between vocals and accompaniment in a song, jointly predicting them with natural expressiveness is a non-trivial task. To this end, we perform extensive explorations into output token patterns, yielding valuable insights. Specifically, (1) in mixed mode, while the model generates high-quality accompaniment, it struggles with natural-sounding vocals. Accompaniment, with higher energy and stable spectral distribution, is easier to produce, whereas vocals, with higher semantic density and a lower signal-to-noise ratio due to overlap, pose a greater challenge. This learning bias makes it difficult to generate vocals with clear lyrics, a problem typically addressed by decoupling and multi-stage methods. To mitigate this issue, we introduce an auxiliary vocal token prediction target, enhancing the model’s focus on vocal features and significantly improving vocal clarity in mixed-token outputs. (2) In dual-track mode, vocals and accompaniment are treated as distinct yet interconnected sequences, generated in sync by a single transformer decoder. We explore various track combination patterns to maintain precise frame-level alignment. Experimental results indicate that the optimal pattern yields well-coordinated vocals and accompaniment, achieving quality on par with mixed-mode generation.
@@
To evaluate the effectiveness of the proposed SongGen framework, we conduct extensive experiments on the MusicCaps (Agostinelli et al., 2023) test set, using both objective and subjective evaluations. The results demonstrate that SongGen generates songs with excellent musicality and vocal-instrument harmony, achieving performance that is competitive with the ground truth. Surprisingly, the generated songs feature expressive vocal techniques, such as vibrato, enhancing naturalness and authenticity. Our contributions can be summarized as follows:
@@ 2.1 Text-to-Music Generation
In recent years, significant progress has been made in text-to-music generation models, which use descriptive text as a condition for controllable music generation. Several works (Agostinelli et al., 2023; Copet et al., 2024) employ transformer-based language models (LMs) (Vaswani et al., 2017) to model sequences of discrete tokens derived from audio codecs (Défossez et al., 2022; Zeghidour et al., 2021; Yang et al., 2023; Kumar et al., 2024). Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Kingma et al., 2021), another competitive class of generative models, have also attained impressive results in music generation(Forsgren & Martiros, 2022; Chen et al., 2024a; Evans et al., 2024; Schneider et al., 2023; Huang et al., 2023; Liu et al., 2024). However, although all the models discussed above excel at generating high-quality instrumental music, they face significant challenges in producing realistic vocals.
@@ 2.2 Song Generation
Recently, a few studies have begun exploring song generation, a task that involves vocal composition, instrumental arrangement, and harmonious generation. One of the pioneering efforts is Jukebox (Dhariwal et al., 2020), which employs a multi-scale VQ-VAE to compress audio into discrete codes and models them using a cascade of transformer models. However, Jukebox offers limited style control, relying solely on genre tags and artist names, and suffers from long inference times. Recently, models like Melodist (Hong et al., 2024) and MelodyLM (Li et al., 2024a) have adopted multi-stage approaches to address the challenges of text-to-song generation. Melodist integrates singing voice synthesis with vocal-to-accompaniment (V2A) techniques, while MelodyLM improves upon Melodist by overcoming its reliance on music scores through a three-stage process: text-to-MIDI, text-to-vocal, and V2A. However, both approaches result in cumbersome training and inference procedures, and their corpus is limited to Mandarin pop songs, lacking diversity. Another model, Song Creator (Lei et al., 2024), utilizes a dual-sequence language model to capture the relationship between vocals and accompaniment. However, it lacks text-based control and produces vocals with limited clarity. Freestyle (Ning et al., 2024) focuses on generating rapping vocals from lyrics and accompaniment inputs but is constrained to a single musical style, with rap typically featuring simpler melodies. Although industry tools like Suno [^1] https://suno.com/ and Udio [^2] https://www.udio.com/ have recently emerged for song generation, neither has disclosed their methodologies or expanded into broader controllable generation tasks. SeedMusic (Bai et al., 2024) leverages both auto-regressive language modeling and diffusion approaches to support song generation. However, SeedMusic is not open-source and relies on a large proprietary dataset, making a fair comparison with our fully open model unfeasible.
@@ 4.1 Experimental setup
Table 1: Objective and Subjective evaluation of Text-to-Song generation. * denotes that we finetune Parlet-tts using our training data. The overall first and second results are marked with bold and underline, respectively. The top subjective results in both of our generation modes are highlighted in yellow.
Model FAD $\downarrow$ KL $\downarrow$ CLAP $\uparrow$ PER $\downarrow$ SECS $\uparrow$ OVL.$\uparrow$ REL.$\uparrow$ VQ.$\uparrow$ HAM. $\uparrow$ SS. $\uparrow$
Ground Truth - - - - - 4.57 4.49 4.49 4.47 4.58
Suno - - - - - 4.28 3.31 4.22 4.33 -
Stable Audio Open (Evans et al., 2024) 4.87 1.15 0.28 - - 3.01 2.87 1.29 - -
MusicGen (Copet et al., 2024) 5.17 0.89 0.09 - - 3.15 2.44 - - -
Parler-tts* (Lyth & King, 2024) 4.13 1.00 0.19 58.61 64.37 2.58 2.13 2.28 2.35 -
Mixed Mixed 1.74 0.71 0.35 51.84 73.69 3.58 3.70 3.55 3.39 3.92
Mixed pro (ours) 1.71 0.69 0.35 40.58 73.78 3.96 3.86 4.07 4.01 4.04
Dual-track Parallel (standard) 2.45 0.75 0.33 48.40 72.27 3.19 3.27 3.36 2.98 3.44
Parallel (V-A) 2.54 0.73 0.33 46.30 72.43 3.36 3.32 3.48 3.08 3.47
Parallel (A-V) 2.31 0.72 0.34 47.00 72.50 3.40 3.33 3.51 3.21 3.51
Interleaving (V-A) 1.96 0.71 0.34 41.82 73.12 3.77 3.69 3.98 3.65 3.88
Interleaving (A-V) (ours) 1.87 0.69 0.35 39.46 73.16 3.95 3.87 4.15 3.82 3.93
@@
Baselines. To the best of our knowledge, no open-source text-to-song model is currently available. We use two state-of-the-art text-to-music models as baselines: Stable Audio Open (Evans et al., 2024) and MusicGen (Copet et al., 2024), both of which generate instrumental music from text. Additionally, we fine-tune Parler-tts (Lyth & King, 2024), a text-to-speech model that generates speech from both transcript and description texts, using our own training data. We also compare our model with Suno, a commercial product, using subjective evaluations.
@@
Evaluation dataset and metrics. For the evaluation dataset, we filter the English-voiced song samples from MusicCaps benchmark (Agostinelli et al., 2023), yielding a test set of 326 samples, with the lyrics annotated by our preprocessing pipeline. We conduct both objective and subjective evaluations. For objective evaluations, Frechet Audio Distance (FAD) measures the generation fidelity; Kullback-Leibler Divergence (KL) evaluates conceptual similarity with the target audio; CLAP Score measures the alignment between the audio and the text description; Speaker Embedding Cosine Similarity (SECS) assesses the similarity of speaker identity; Phoneme Error Rate (PER) gauges adherence to the provided lyrics. Note that due to limitations in the ASR model, the PER values are higher than the actual error rate, but the relative differences remain meaningful. For each method, we generate the audio five times with different random seeds and report the average metric. For subjective evaluations, we employ Mean Opinion Score (MOS) tests, assessing five key aspects: overall quality (OVL.), focusing on musicality and naturalness; relevance to the text description (REL.); vocal quality, with an emphasis on clarity and intelligibility (VQ.); harmony between vocals and accompaniment (HAM.); and similarity to the original singer (SS.). The appendix C shows details of the evaluations.
@@ 4.2 Results of Text-to-song Generation
As shown in Table 1, we compare our models, including mixed mode and dual-track mode, with several baselines. For all our models in this table, we use the first 3-second vocal clip of the ground truth as the reference voice.
Comparison with Baselines. SongGen significantly outperforms Stable Audio Open, MusicGen, and Parler-tts across both subjective and objective metrics. The test set contains many voice-related descriptions, such as “A group of female vocalists sings this energetic swing song.” However, MusicGen generates pure music, lacking any vocal signals, which results in exceptionally low CLAP scores. Stable Audio Open generates some vocal signals based on the input text, but these signals do not form recognizable words. Although Parler-tts has achieved remarkable success in controllable text-to-speech tasks, fine-tuning it for the text-to-song task proves to be ineffective. This highlights the greater complexity of the text-to-song task compared to text-to-speech.
Although SongGen shows some gaps when compared to Ground Truth and Suno, it is important to note that we use only 2k hours of labeled data, sourced from publicly available datasets. Despite the limited data, SongGen achieves competitive performance on an out-of-domain test dataset. Figure 4 shows a mel-spectrogram of our generated songs, demonstrating that SongGen produces songs with various singing techniques like vibrato. Compared to Suno, a commercial product, SongGen outperforms in terms of text relevance and vocal control. Suno struggles to adhere to the highly detailed textual descriptions in MusicCaps (as shown by the REL. metric) and lacks voice cloning support, which gives our model a clear advantage in these aspects.

### v2

@@ Abstract
Text-to-song generation, the task of creating vocals and accompaniment from textual inputs, poses significant challenges due to domain complexity and data scarcity. Existing approaches often employ multi-stage generation procedures, leading to cumbersome training and inference pipelines, as well as suboptimal overall generation quality due to error accumulation across stages. In this paper, we propose SongGen, a fully open-source, single-stage auto-regressive transformer designed for controllable song generation. The proposed model facilitates fine-grained control over diverse musical attributes, including lyrics and textual descriptions of instrumentation, genre, mood, and timbre, while also offering an optional three-second reference clip for voice cloning. Within a unified auto-regressive framework, SongGen supports two output modes: mixed mode, which generates a mixture of vocals and accompaniment directly, and dual-track mode, which synthesizes them separately for greater flexibility in downstream applications. We explore diverse token pattern strategies for each mode, leading to notable improvements and valuable insights. Furthermore, we design an automated data preprocessing pipeline with effective quality control. To foster community engagement and future research, we will release our model weights, training code, annotated data, and preprocessing pipeline. The code is available at https://github.com/LiuZH-19/SongGen.
@@ 1 Introduction
Figure 1: Multi-stage approaches generate vocals and accompaniment sequentially, leading to cumbersome pipelines and cumulative errors. SongGen simplifies this with a single-stage auto-regressive transformer that jointly models both tracks, supporting mixed and dual-track modes with better quality and efficiency.
Song generation presents greater complexity than speech or instrumental music generation (Lyth & King, 2024; Chen et al., 2024b; Liu et al., 2024; Copet et al., 2024). Unlike speech, singing spans a broader pitch range, incorporates a wider variety of expressive vocal techniques, and follows more dynamic rhythmic patterns. Moreover, song generation requires precise alignment between vocal and instrumental components to ensure musicality, harmonic consistency, and structural coherence—making it a uniquely challenging and underexplored problem. The scarcity of open-source data further limits research in this area.
Recent text-to-song approaches (Hong et al., 2024; Li et al., 2024a) decompose songs into separate vocal and accompaniment tracks and adopt multi-stage generation pipelines. As illustrated in Figure 1, these models first generate the vocal track from lyrics, then produce the accompaniment using natural language prompts alongside the generated vocals. However, such pipeline-based approach often fail to capture global optimality due to error accumulation across stages. This limitation is especially problematic for song generation. For instance, in genres like rap, vocal rhythm is tightly coupled with the instrumental beat. Generating vocals first without considering the underlying rhythm may result in rhythm misalignment. Conversely, in expressive genres such as ballads, where vocals typically guide the emotional flow, generating accompaniment first may constrain vocal expressiveness, resulting in rigid or disconnected performances. In both cases, pipeline approaches struggle to capture the intricate interplay between vocals and accompaniment. In addition, multi-stage generation results in cumbersome training and inference pipelines. Given these limitations in generation quality and efficiency, an important question arises: Is it possible for a single-stage model to achieve effective text-to-song generation?
@@
However, due to the sophisticated relationship between vocals and accompaniment in a song, jointly predicting them with natural expressiveness is a non-trivial task. To this end, we perform extensive explorations into output token patterns, yielding valuable insights. Specifically, (1) in mixed mode, while the model generates high-quality accompaniment, it struggles with natural-sounding vocals. Accompaniment, with higher energy and stable spectral distribution, tends to converge faster during training, whereas vocals, with higher semantic density and a lower signal-to-noise ratio due to overlap, present greater modeling challenges. This learning bias makes it difficult to generate vocals with clear lyrics, a problem typically addressed by decoupling and multi-stage methods. To mitigate this issue, we introduce an auxiliary vocal token prediction target, enhancing the model’s focus on vocal features and significantly improving vocal clarity in mixed-token outputs. (2) In dual-track mode, vocals and accompaniment are treated as distinct yet interconnected sequences, generated in sync by a single transformer decoder. We explore various track combination patterns to maintain precise frame-level alignment. Experimental results indicate that the optimal pattern yields well-coordinated vocals and accompaniment, achieving quality on par with mixed-mode generation.
@@
To evaluate the effectiveness of the proposed SongGen framework, we conduct extensive experiments on the MusicCaps (Agostinelli et al., 2023) test set. The results demonstrate that SongGen outperforms the multi-stage baseline and generates songs with excellent musicality and vocal-instrument harmony, achieving performance that is competitive with the ground truth. Surprisingly, the generated songs feature expressive vocal techniques, such as vibrato, enhancing naturalness and authenticity. Our contributions can be summarized as follows:
@@ 2.1 Text-to-Music Generation
In recent years, significant progress has been made in text-to-music generation models, which use descriptive text as a condition for controllable music generation. Several works (Agostinelli et al., 2023; Copet et al., 2024) employ transformer-based language models (LMs) (Vaswani et al., 2017) to model sequences of discrete tokens derived from audio codecs (Défossez et al., 2022; Zeghidour et al., 2021; Yang et al., 2023; Kumar et al., 2024). Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Kingma et al., 2021), another competitive class of generative models, have also attained impressive results in music generation(Forsgren & Martiros, 2022; Chen et al., 2024a; Evans et al., 2024; Schneider et al., 2023; Huang et al., 2023; Liu et al., 2024). While most models generate a mixture of stems, MusicGen-Stem (Rouard et al., 2025) proposes a multi-stem generative model with three separate tracks (bass, drums and other), suggesting that disentangling components can facilitate generation and source editing. However, although all the models discussed above excel at generating high-quality instrumental music, they face significant challenges in producing realistic vocals.
@@ 2.2 Song Generation
Recently, a few studies have begun exploring song generation, a task that involves vocal composition, instrumental arrangement, and harmonious generation. One of the pioneering efforts is Jukebox (Dhariwal et al., 2020), which employs a multi-scale VQ-VAE to compress audio into discrete codes and models them using a cascade of transformer models. However, Jukebox offers limited style control, relying solely on genre tags and artist names, and suffers from long inference times. Recently, models like Melodist (Hong et al., 2024) and MelodyLM (Li et al., 2024a) have adopted multi-stage approaches to address the challenges of text-to-song generation. Melodist integrates singing voice synthesis with vocal-to-accompaniment (V2A) techniques, while MelodyLM improves upon Melodist by overcoming its reliance on music scores through a three-stage process: text-to-MIDI, text-to-vocal, and V2A. However, both approaches result in cumbersome training and inference procedures, and their corpus is limited to Mandarin pop songs, lacking diversity. Another model, SongCreator (Lei et al., 2024), utilizes a dual-sequence language model to capture the relationship between vocals and accompaniment. However, it lacks text-based control and produces vocals with limited clarity. Freestyle (Ning et al., 2024) focuses on generating rapping vocals from lyrics and accompaniment inputs but is constrained to a single musical style, with rap typically featuring simpler melodies. The concurrent work Yue (Yuan et al., 2025) achieves impressive results by scaling up both data and model size, adopting a track-decoupled next-token prediction and a two-stage causal LM framework. In contrast, we explore more diverse and efficient token pattern designs within a single-stage transformer. Although industry tools like Suno [^1] https://suno.com/ and Udio [^2] https://www.udio.com/ have recently emerged for song generation, neither has disclosed their methodologies or expanded into broader controllable generation tasks. SeedMusic (Bai et al., 2024) leverages both auto-regressive language modeling and diffusion approaches to support song generation. However, SeedMusic is not open-source and relies on a large proprietary dataset, making a fair comparison with our fully open model unfeasible.
@@ 4.1 Experimental setup
Table 1: Automatic evaluation of Text-to-Song generation. * denotes that we finetune Parlet-tts using our training data. The top two results, excluding the ground truth, are marked in bold and underlined, respectively.
Model Metric Distrib. Match Alignment Aesthetics
[…]
Ground Truth - - 0.18 0.105 21.39 76.42 7.08 7.04 6.41 7.30
Parler-tts* 4.13 1.00 0.19 0.074 58.61 64.37 5.96 6.62 5.49 6.82
Multi-Stage 2.18 0.78 0.29 0.085 38.80 74.04 6.39 6.27 5.90 6.69
Mixed Mixed 1.74 0.71 0.35 0.093 51.84 73.69 6.50 6.66 6.14 7.03
Mixed pro (ours) 1.71 0.69 0.35 0.094 40.58 73.78 6.77 6.86 6.18 7.19
Dual-track Parallel (standard) 2.45 0.75 0.33 0.087 48.40 72.27 6.21 6.40 5.68 6.88
Parallel (V-A) 2.54 0.73 0.33 0.088 46.30 72.43 6.21 6.49 5.72 6.92
Parallel (A-V) 2.31 0.72 0.34 0.089 47.00 72.50 6.26 6.47 5.80 6.92
Interleaving (V-A) 1.96 0.71 0.34 0.092 41.82 73.12 6.52 6.52 5.97 7.03
Interleaving (A-V) (ours) 1.87 0.69 0.35 0.093 39.46 73.16 6.67 6.72 6.11 7.12
@@
Baselines. To the best of our knowledge, no open-source text-to-song model is currently available. We therefore conduct a controlled comparison between single-stage and multi-stage approaches. The multi-stage baseline consists of two transformer models with the same architecture and training data as SongGen. Specifically, in Stage 1, the first model generates the vocal track from the lyrics, description, and a 3-second reference voice. In Stage 2, the second model generates the accompaniment, conditioned on the same textual inputs and the generated vocal (prepended into the decoder). The final song is then mixed from the two tracks. Additionally, we fine-tune Parler-tts (Lyth & King, 2024), a text-to-speech model that generates speech from both transcript and description texts, using our own training data. We also compare our model with Suno, a commercial product, using human evaluations.
Evaluation dataset and metrics. For the evaluation dataset, we filter the English-voiced song samples from MusicCaps benchmark (Agostinelli et al., 2023), yielding a test set of 326 samples, with the lyrics annotated by our preprocessing pipeline.
@@
For automatic evaluations, Frechet Audio Distance (FAD) measures the generation fidelity; Kullback-Leibler Divergence (KL) evaluates conceptual similarity with the target audio; CLAP Score and CLaMP3 score (Wu et al., 2025) measures the alignment between the audio and the text description; Speaker Embedding Cosine Similarity (SECS) assesses the similarity of speaker identity; Phoneme Error Rate (PER) gauges adherence to the provided lyrics. Note that due to limitations in the ASR model, the PER values are higher than the actual error rate, but the relative differences remain meaningful. We also introduce content-based aesthetics metrics (Tjandra et al., 2025), covering Content Enjoyment (CE), Content Usefulness (CU), Production Complexity (PC), and Production Quality (PQ). For each method, we generate the audio five times with different random seeds and report the average metric.
For human evaluations, we employ Mean Opinion Score (MOS) tests, assessing five key aspects: overall quality (OVL.), focusing on musicality and naturalness; relevance to the text description (REL.); vocal quality, with an emphasis on clarity and intelligibility of the singing voice (VQ.); harmony between vocals and accompaniment (HAM.); and similarity to the original singer (SS.). The appendix B shows details of the evaluations.
@@ 4.2 Results of Text-to-song Generation
Tables 1 and 2 report the automatic and human evaluation results, respectively, for our mixed and dual-track models alongside several baselines. In both tables, the first 3-second vocal clip from the ground-truth is used as the reference voice for all our models and the multi-stage baseline.
Table 2: Humain evaluation of Text-to-Song generation. The overall first and second results are marked with bold and underline, respectively. The top results in both of our generation modes are highlighted in yellow.
Model OVL.$\uparrow$ REL.$\uparrow$ VQ.$\uparrow$ HAM. $\uparrow$ SS. $\uparrow$
Ground Truth 4.57 $\pm$0.04 4.49 $\pm$0.03 4.49 $\pm$0.05 4.47 $\pm$0.04 4.58 $\pm$0.03
Suno 4.28$\pm$0.04 3.31$\pm$ 0.04 4.22$\pm$ 0.05 4.33$\pm$ 0.05 -
Parler-tts* 2.58$\pm$ 0.06 2.13$\pm$0.05 2.28$\pm$0.03 2.35$\pm$0.04 -
Multi-Stage 3.39$\pm$0.03 3.20$\pm$0.04 3.98$\pm$0.07 2.97$\pm$0.04 3.89$\pm$0.03
Mixed Mixed 3.58$\pm$0.05 3.70$\pm$0.02 3.55$\pm$ 0.07 3.39$\pm$ 0.05 3.92$\pm$0.05
Mixed pro 3.96 $\pm$0.04 3.86$\pm$ 0.04 4.07$\pm$0.06 4.01$\pm$0.05 4.04$\pm$0.05
Dual-track Parallel (std.) 3.19$\pm$0.04 3.27$\pm$0.06 3.36$\pm$0.04 2.98$\pm$0.05 3.44$\pm$0.03
Parallel (V-A) 3.36$\pm$0.03 3.32$\pm$0.05 3.48$\pm$0.05 3.08$\pm$0.06 3.47$\pm$0.04
Parallel (A-V) 3.40$\pm$0.03 3.33$\pm$0.04 3.51$\pm$0.04 3.21$\pm$0.05 3.51$\pm$0.05
Inter. (V-A) 3.77$\pm$0.03 3.69$\pm$0.05 3.98$\pm$0.06 3.65$\pm$0.04 3.88$\pm$0.04
Inter. (A-V) 3.95$\pm$0.03 3.87$\pm$0.06 4.15$\pm$0.05 3.82$\pm$ 0.03 3.93$\pm$0.04
Comparison with baselines. Although Parler-TTS excels in controllable text-to-speech, fine-tuning it for text-to-song is ineffective, as shown in the tables. This highlights the greater complexity of generating expressive vocals and musically coherent accompaniment in text-to-song tasks, compared to conventional speech synthesis.
Compared to the multi-stage baseline, our single-stage model outperforms it across both automatic and human evaluations. The only exceptions are the PER and SECS metrics, where the multi-stage model performs slightly better—due to its first-stage training being focused exclusively on clean vocal targets. Nonetheless, our single-stage approach shows clear strengths on all other metrics, particularly in aesthetics-related aspects: CE (+5.9%), CU (+9.4%), PC (+4.7%), and PQ (+7.5%), as well as in human evaluations, with Overall Quality (OVL.) increasing by 0.57 and Harmony (HAM.) by 1.04 on a five-point MOS scale. By directly modeling the joint distribution $P(\text{vocal},\text{accompaniment})$, rather than optimizing $P(\text{vocal})$ and $P(\text{accompaniment}\mid\text{vocal})$ separately, the single-stage approach facilitates more effective global coordination and avoids error accumulation across stages. This is particularly beneficial in song generation, where the alignment and coherence between vocals and accompaniment are essential. We also showcase generation samples from both the multi-stage and single-stage models on our demo page  [^3] https://liuzh-19.github.io/SongGen/, where the multi-stage outputs are often perceptibly off-beat. In addition to better quality, the single-stage model is also more efficient. On an A800 GPU, our mixed-pro model averages 18.04 seconds to generate a 30-second sample, compared to 42.85 seconds for the multi-stage baseline.
While SongGen shows some gaps when compared to Ground Truth and Suno, it is worth noting that we use only 2k hours of labeled data, sourced from publicly available datasets. Despite the limited data, SongGen achieves competitive performance. Figure 4 shows a mel-spectrogram of our generated songs, demonstrating that SongGen produces songs with various singing techniques like vibrato. Compared to Suno, a commercial product, SongGen outperforms in terms of text relevance and vocal control. Suno struggles to adhere to the highly detailed textual descriptions in MusicCaps (as shown by the REL. metric) and lacks voice cloning support, giving our model a distinct advantage in these aspects.